# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v2

@@ 6 Conclusion
Experimental code related to this paper is provided in the Supplementary Material. Detailed theoretical derivations are provided in App. A. The complete architecture of HOG-Diff, along with the training objectives and sampling procedures, is described in App. C. Details of the datasets, preprocessing steps, and experimental settings are provided in App. E. Additional experimental results, ablation studies, and visualizations can be found in App. F.
In this work, Large Language Models (LLMs) are used solely for language polishing.
Acknowledgments
@@ A.1 Diffusion Bridge Process
The Doob’s $h$-transform can modify an SDE to pass through a specified endpoint (Doob & Doob, 1984). When applied to the GOU process, it eliminates variance in the terminal state by driving the diffusion process toward a Dirac distribution centered at $\bm{G}_{\tau_{k}}$ (Heng et al., 2021; Yue et al., 2024), making it well-suited for stochastic modelling with terminal constraints.
In the following, we derive the generalized Ornstein–Uhlenbeck (GOU) bridge process using Doob’s $h$-transform (Doob & Doob, 1984), and subsequently examine its relationship with the Brownian bridge process.
@@
Let $\bm{G}_{t}$ evolve according to the generalized OU process in Eq. 4, subject to the terminal conditional $\bm{\mu}=\bm{G}_{\tau_{k}}$. Applying Doob’s $h$-transform (Doob & Doob, 1984), we can derive the GOU bridge process as follows:
@@
$\begin{split}\bar{\mathbf{m}}_{t}=\mathbb{E}[\mathbf{x}_{t}\mid\mathbf{x}_{0}\mid\mathbf{x}_{T}]=\mathbb{E}[\mathbf{x}_{t}\mid\mathbf{x}_{0}]+\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})\mathrm{Var}(\mathbf{x}_{T}\mid\mathbf{x}_{0})^{-1}(\mathbf{x}_{T}-\mathbb{E}[\mathbf{x}_{T}\mid\mathbf{x}_{0}]),\\ \bar{v}_{t}^{2}=\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0}\mid\mathbf{x}_{T})=\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})\mathrm{Var}(\mathbf{x}_{T}\mid\mathbf{x}_{0})^{-1}\mathrm{Cov}(\mathbf{x}_{T},\mathbf{x}_{t}\mid\mathbf{x}_{0}).\end{split}$ (19)
@@
$\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})=\mathrm{Cov}\left(\mathbf{x}_{t},(\mathbf{x}_{t}-\mathbf{x}_{T})e^{-\bar{\theta}_{t:T}}\mid\mathbf{x}_{0}\right)=e^{-\bar{\theta}t:T}\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0}).$ (20)
@@
$\begin{split}\bar{v}_{t}^{2}&=v_{t}^{2}-\left(e^{-\bar{\theta}_{t:T}}v_{t}^{2}\right)^{2}/v_{T}^{2}\\ &=\frac{v_{t}^{2}}{v_{T}^{2}}(v_{T}^{2}-e^{-2\bar{\theta}_{t:T}}v_{t}^{2})\\ &=\frac{v_{t}^{2}}{v_{T}^{2}}\sigma^{2}\left(1-e^{-2\bar{\theta}_{T}}-e^{-2\bar{\theta}_{t:T}}(1-e^{-\bar{2\theta}_{t}})\right)\\ &=v_{t}^{2}v_{t:T}^{2}/v_{T}^{2}.\end{split}$ (22)
@@ A.2 Proof of Theorem 3
To establish proof, we begin by introducing essential definitions and assumptions.
@@
Let $\ell^{(k)}(\bm{\theta})$ be a loss function that is $\beta$-smooth and satisfies the $\mu$-PL (Polyak-Łojasiewicz) condition in the ball $B\left(\bm{\theta}_{0},R\right)$ of radius $R=2N\sqrt{2\beta\ell^{(k)}\left(\bm{\theta}_{0}\right)}/(\mu\delta)$, where $\delta>0$. Then, with probability $1-\delta$ over the choice of mini-batch of size $b$, stochastic gradient descent (SGD) with a learning rate $\eta^{*}=\frac{\mu N}{N\beta\left(N^{2}\beta+\mu(b-1)\right)}$ converges to a global solution in the ball $B$ with exponential convergence rage:
@@
Consider the loss function of classical generative models goes as: $\bar{\ell}(\bm{\varphi})=\mathbb{E}||\bm{s}_{\bm{\varphi}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}$. To demonstrate that the proposed framework converges faster to the target distribution compared to the classical one-step generation framework, it suffices to show that: $\beta_{\nabla p_{t|\tau_{k-1}}}\leq\beta_{\nabla q_{t|0}}$ and $\beta_{\nabla^{2}p_{t|\tau_{k-1}}}\leq\beta_{\nabla^{2}q_{t|0}}$.
Let $\mathbf{x}\sim q_{t|0}$ and $\mathbf{x}^{\prime}\sim p_{t|\tau_{k-1}}$. Since we inject topological information from $\mathbf{x}$ into $\mathbf{x}^{\prime}$, $\mathbf{x}^{\prime}$ can be viewed as being obtained by adding noise to $\mathbf{x}$. Hence, we can model $\mathbf{x}^{\prime}$ as $\mathbf{x}^{\prime}=\mathbf{x}+\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})$. The variance of Gaussian noise $\sigma^{2}$ controls the information remained in $z^{\prime}$. Hence, its distribution can be expressed as $p(\mathbf{x}^{\prime})=\int q(\mathbf{x}^{\prime}-\epsilon)\pi(\epsilon)\mathop{}\!\mathrm{d}\epsilon$.
@@ A.3 Proof of Theorem 4
Under appropriate Lipschitz and boundedness assumptions, the reconstruction error of HOG-Diff satisfies the following bound:
$\mathcal{E}(0)\leq\alpha(0)\exp{\int_{0}^{\tau_{1}}\gamma(s)}\mathop{}\!\mathrm{d}{s},$ (37)
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than that of classical graph generation models.
@@
Let $\mathcal{E}(t)=\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$, which reflects the expected error between the data reconstructed with the ground truth score $\nabla\log p_{t}(\cdot)$ and the learned scores $\bm{s}_{\bm{\theta}}(\cdot)$. In particular, $\bar{\bm{G}}$ is obtained by solving the following oracle reversed time SDE:
@@
Here, $\mathbf{f}_{k,t}$ is the drift function of the Ornstein–Uhlenbeck bridge. For simplicity, we we denote the Lipschitz norm by $||\cdot||_{\operatorname{lip}}$ and $\mathbf{f}_{k,s}(\bm{G}_{s})=h_{k,s}(\bm{G}_{\tau_{k}}-\bm{G}_{s})$, where $h_{k,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{k}}}-1}\right)$.
To bound the expected reconstruction error $\mathbb{E}\left\|\bar{\bm{G}}_{\tau_{k-1}}-\widehat{\bm{G}}_{\tau_{k-1}}\right\|^{2}$ at each generation process, we begin by analyze how $\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$ evolves as time $t$ is reversed from $\tau_{k}$ to $\tau_{k-1}$. The reconstruction error goes as follows
@@
$\leq\underbrace{C^{2}\ell^{(k)}(\bm{\theta})\int_{\tau_{k}}^{t}g_{k,s}^{4}\mathrm{d}\bar{s}+C\mathcal{E}(\tau_{k})\int_{\tau_{k}}^{t}h_{k,s}^{2}\mathrm{d}\bar{s}}_{\alpha(t)}+\int_{\tau_{k}}^{t}\underbrace{\left(C^{2}g_{k,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{k,s}\|_{\mathrm{lip}}^{2}\right)}_{\gamma(s)}\mathcal{E}(s)\mathrm{d}\bar{s}$
$=\alpha(t)+\int_{\tau_{k}}^{t}\gamma(s)\mathcal{E}(s)\mathrm{d}\bar{s}.$
@@
$v(t)=\mathcal{E}(\tau_{k}-t)\leq\alpha(\tau_{k}-t)+\int_{0}^{t}\gamma(\tau_{k}-s^{\prime})v(s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}.$ (41)
Here, $\alpha(\tau_{k}-t)$ is a non-decreasing function. By applying Grönwall’s inequality, we can derive that
$v(t)$ $\leq\alpha(\tau_{k}-t)\exp{\int_{0}^{t}\gamma(\tau_{k}-s^{\prime})}\mathrm{d}s^{\prime}$ (42)
$=\alpha(\tau_{k}-t)\exp{\int_{\tau_{k}-t}^{\tau_{k}}\gamma(s)}\mathrm{d}s.$ (43)

### v3

@@ Acknowledgments
Detailed theoretical derivations are provided in App. A. The complete architecture of HOG-Diff, along with the training and sampling procedures, is described in App. C. Details of the datasets, preprocessing steps, and experimental settings are provided in App. E. Additional experimental results, ablation studies, and visualizations can be found in App. F.
@@ A.1 Diffusion Bridge Process
Doob’s $h$-transform can modify an SDE to pass through a specified endpoint (Doob and Doob, 1984). When applied to the GOU process, it eliminates variance in the terminal state by driving the diffusion process toward a Dirac distribution centered at $\bm{G}_{\tau_{k}}$ (Heng et al., 2021; Yue et al., 2024), making it well-suited for stochastic modelling with terminal constraints.
In the following, we derive the generalized Ornstein–Uhlenbeck (GOU) bridge process using Doob’s $h$-transform (Doob and Doob, 1984), and subsequently examine its relationship with the Brownian bridge process.
@@
Let $\bm{G}_{t}$ evolve according to the generalized OU process in Eq. 5, subject to the terminal conditional $\bm{\mu}=\bm{G}_{\tau_{k}}$. Applying Doob’s $h$-transform (Doob and Doob, 1984), we can derive the GOU bridge process as follows:
@@
$\begin{split}\bar{\mathbf{m}}_{t}=\mathbb{E}[\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T}]=\mathbb{E}[\mathbf{x}_{t}\mid\mathbf{x}_{0}]+\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})\mathrm{Var}(\mathbf{x}_{T}\mid\mathbf{x}_{0})^{-1}(\mathbf{x}_{T}-\mathbb{E}[\mathbf{x}_{T}\mid\mathbf{x}_{0}]),\\ \bar{v}_{t}^{2}=\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})\mathrm{Var}(\mathbf{x}_{T}\mid\mathbf{x}_{0})^{-1}\mathrm{Cov}(\mathbf{x}_{T},\mathbf{x}_{t}\mid\mathbf{x}_{0}).\end{split}$ (19)
@@
$\mathrm{Cov}(\mathbf{x}_{t},\mathbf{x}_{T}\mid\mathbf{x}_{0})=\mathrm{Cov}\left(\mathbf{x}_{t},(\mathbf{x}_{t}-\mathbf{x}_{T})e^{-\bar{\theta}_{t:T}}\mid\mathbf{x}_{0}\right)=e^{-\bar{\theta}_{t:T}}\mathrm{Var}(\mathbf{x}_{t}\mid\mathbf{x}_{0}).$ (20)
@@
$\begin{split}\bar{v}_{t}^{2}&=v_{t}^{2}-\left(e^{-\bar{\theta}_{t:T}}v_{t}^{2}\right)^{2}/v_{T}^{2}\\ &=\frac{v_{t}^{2}}{v_{T}^{2}}(v_{T}^{2}-e^{-2\bar{\theta}_{t:T}}v_{t}^{2})\\ &=\frac{v_{t}^{2}}{v_{T}^{2}}\sigma^{2}\left(1-e^{-2\bar{\theta}_{T}}-e^{-2\bar{\theta}_{t:T}}(1-e^{-2\bar{\theta}_{t}})\right)\\ &=v_{t}^{2}v_{t:T}^{2}/v_{T}^{2}.\end{split}$ (22)
@@ A.2 Proof of Theorem 3
To establish the proof, we begin by introducing essential definitions and assumptions.
@@
Let $\ell^{(k)}(\bm{\theta})$ be a loss function that is $\beta$-smooth and satisfies the $\mu$-PL (Polyak-Łojasiewicz) condition in the ball $B\left(\bm{\theta}_{0},R\right)$ of radius $R=2N\sqrt{2\beta\ell^{(k)}\left(\bm{\theta}_{0}\right)}/(\mu\delta)$, where $\delta>0$. Then, with probability $1-\delta$ over the choice of mini-batch of size $b$, stochastic gradient descent (SGD) with a learning rate $\eta^{*}=\frac{\mu N}{N\beta\left(N^{2}\beta+\mu(b-1)\right)}$ converges to a global solution in the ball $B$ with exponential convergence rate:
@@
The loss function used in classical diffusion generative models is given by $\bar{\ell}(\bm{\varphi})=\mathbb{E}||\bm{s}_{\bm{\varphi}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}$. To demonstrate that the proposed framework converges faster to the target distribution compared to the classical one-step generation framework, it suffices to show that: $\beta_{\nabla p_{t|\tau_{k-1}}}\leq\beta_{\nabla q_{t|0}}$ and $\beta_{\nabla^{2}p_{t|\tau_{k-1}}}\leq\beta_{\nabla^{2}q_{t|0}}$.
Let $\mathbf{x}\sim q_{t|0}$ and $\mathbf{x}^{\prime}\sim p_{t|\tau_{k-1}}$. Since we inject topological information from $\mathbf{x}$ into $\mathbf{x}^{\prime}$, $\mathbf{x}^{\prime}$ can be viewed as being obtained by adding noise to $\mathbf{x}$. Hence, we can model $\mathbf{x}^{\prime}$ as $\mathbf{x}^{\prime}=\mathbf{x}+\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})$. The variance of Gaussian noise $\sigma^{2}$ controls the information retained in $\mathbf{x}^{\prime}$. Hence, its distribution can be expressed as $p(\mathbf{x}^{\prime})=\int q(\mathbf{x}^{\prime}-\epsilon)\pi(\epsilon)\mathop{}\!\mathrm{d}\epsilon$.
@@ A.3 Proof of Theorem 4
Before comparing the reconstruction error bounds of HOG-Diff and the classical diffusion model, we first relate their optimal score-matching losses at the function-class level.
Lemma 6.
Let $\ell^{(k)}(\bm{\theta}^{(k)})$ denote the score-matching loss of the $k$-th HOG-Diff stage on $\tau_{k-1},\tau_{k}]$, and let $\ell_{\mathrm{cls}}(\bm{\varphi})$ be the classical single-stage loss on $[0,T]$. Define the corresponding optimal values
$\ell^{(k)}_{\star}:=\inf_{\bm{\theta}^{(k)}}\ell^{(k)}(\bm{\theta}^{(k)}),\qquad\ell_{\mathrm{cls},\star}:=\inf_{\bm{\varphi}}\ell_{\mathrm{cls}}(\bm{\varphi}).$ (37)
Assume that, for each $k$, the score network used in HOG-Diff on $\tau_{k-1},\tau_{k}]$ is instantiated from the same backbone architecture as the classical score network on $[0,T]$ (with the coarse graph $\bm{G}_{\tau_{k}}$ provided as an additional input that can be ignored by a suitable parameter choice). Under this construction, the classical score class is contained in each stage-wise score class, and we have
$\sum_{k=1}^{K}\ell^{(k)}_{\star}\;\leq\;\ell_{\mathrm{cls},\star}.$ (38)
Proof.
Fix any parameter $\bm{\varphi}$ of the classical model and consider the associated score network $\bm{s}_{\bm{\varphi}}(\bm{G}_{t},t)$ on $[0,T]$. By the expressivity assumption, for each $k$ there exists a parameter vector $\bm{\theta}^{(k)}$ such that the HOG-Diff score network on $\tau_{k-1},\tau_{k}]$ can represent exactly the same score function while ignoring the coarse guide:
$\bm{s}^{(k)}_{\bm{\theta}^{(k)}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)=\bm{s}_{\bm{\varphi}}(\bm{G}_{t},t),\qquad t\in[\tau_{k-1},\tau_{k}].$ (39)
By definition, the stage-wise loss $\ell^{(k)}(\bm{\theta}^{(k)})$ is obtained by restricting the classical objective $\ell_{\mathrm{cls}}(\bm{\varphi})$ to the time window $\tau_{k-1},\tau_{k}]$ and using the same score function on that interval. Therefore, under the above parameter choice we have
$\ell^{(k)}(\bm{\theta}^{(k)})=\ell_{\mathrm{cls}}^{(k)}(\bm{\varphi}),$ (40)
where $\ell_{\mathrm{cls}}^{(k)}(\bm{\varphi})$ denotes the contribution of the interval $\tau_{k-1},\tau_{k}]$ to the classical loss. Summing over all disjoint windows yields the exact decomposition
$\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta}^{(k)})=\sum_{k=1}^{K}\ell_{\mathrm{cls}}^{(k)}(\bm{\varphi})=\ell_{\mathrm{cls}}(\bm{\varphi}).$ (41)
Taking the infimum over $(\bm{\theta}^{(1)},\dots,\bm{\theta}^{(K)})$ on the left-hand side and over $\bm{\varphi}$ on the right-hand side gives
$\sum_{k=1}^{K}\ell^{(k)}_{\star}=\sum_{k=1}^{K}\inf_{\bm{\theta}^{(k)}}\ell^{(k)}(\bm{\theta}^{(k)})\;\leq\;\inf_{\bm{\varphi}}\ell_{\mathrm{cls}}(\bm{\varphi})=\ell_{\mathrm{cls},\star}.$ (42)
Sec. A.3 reflects the structural fact that the classical model is a special case of the multi-stage architecture when all stages share the same backbone and the coarse graph input is ignored. Intuitively, each HOG-Diff stage only needs to approximate a smoother conditional score (given the coarse graph), whereas the classical model has to fit the full marginal score in a single stage. For a fixed network capacity, decomposing the problem into these easier conditional subproblems is therefore not expected to deteriorate the overall score-matching accuracy.
@@
Under standard Lipschitz and boundedness assumptions, the reconstruction error of HOG-Diff satisfies the following bound:
$\mathcal{E}_{\mathrm{hog}}(0)\leq A_{\mathrm{hog}}\,e^{T_{\mathrm{hog}}},$ (43)
where $A_{\mathrm{hog}}=C^{2}\sum_{k=1}^{K}\ell^{(k)}(\bm{\theta})\int_{\tau_{k-1}}^{\tau_{k}}g_{k,s}^{4}\mathrm{d}s$ and $T_{\mathrm{hog}}=\sum_{k=1}^{K}\int_{\tau_{k-1}}^{\tau_{k}}\gamma_{k}(s)\mathrm{d}s+C\sum_{k=1}^{K-1}\int_{\tau_{k-1}}^{\tau_{k}}h_{k,s}^{2}\mathrm{d}s$. Moreover, under the same assumptions, HOG-Diff admits a tighter reconstruction error bound than classical diffusion models.
@@
Recall that $\mathcal{E}(t)=\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$ reflects the expected error between the data reconstructed with the ground truth score $\nabla\log p_{t}(\cdot)$ and the learned scores $\bm{s}_{\bm{\theta}}(\cdot)$. In particular, $\bar{\bm{G}}$ is obtained by solving the following oracle reversed time SDE:
@@
Here, $\mathbf{f}_{k,t}$ is the drift function of the Ornstein–Uhlenbeck bridge. For simplicity, we denote the Lipschitz norm by $||\cdot||_{\operatorname{lip}}$ and $\mathbf{f}_{k,s}(\bm{G}_{s})=h_{k,s}(\bm{G}_{\tau_{k}}-\bm{G}_{s})$, where $h_{k,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{k}}}-1}\right)$.
Step 1: Single stage error bound.
To bound the expected reconstruction error $\mathbb{E}\left\|\bar{\bm{G}}_{\tau_{k-1}}-\widehat{\bm{G}}_{\tau_{k-1}}\right\|^{2}$ at each generation process, we begin by analyzing how $\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$ evolves as time $t$ is reversed from $\tau_{k}$ to $\tau_{k-1}$. The reconstruction error goes as follows
@@
$\leq\underbrace{C^{2}\ell^{(k)}(\bm{\theta})\int_{\tau_{k}}^{t}g_{k,s}^{4}\mathrm{d}\bar{s}+C\mathcal{E}(\tau_{k})\int_{\tau_{k}}^{t}h_{k,s}^{2}\mathrm{d}\bar{s}}_{\alpha_{k}(t)}+\int_{\tau_{k}}^{t}\underbrace{\left(C^{2}g_{k,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{k,s}\|_{\mathrm{lip}}^{2}\right)}_{\gamma_{k}(s)}\mathcal{E}(s)\mathrm{d}\bar{s}$
$=\alpha_{k}(t)+\int_{\tau_{k}}^{t}\gamma_{k}(s)\mathcal{E}(s)\mathrm{d}\bar{s}.$
@@
$v(t)=\mathcal{E}(\tau_{k}-t)\leq\alpha_{k}(\tau_{k}-t)+\int_{0}^{t}\gamma_{k}(\tau_{k}-s^{\prime})v(s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}.$ (47)
Here, $\alpha_{k}(\tau_{k}-t)$ is a non-decreasing function. By applying Grönwall’s inequality, we can derive that
$v(t)$ $\leq\alpha_{k}(\tau_{k}-t)\exp{\int_{0}^{t}\gamma_{k}(\tau_{k}-s^{\prime})}\mathrm{d}s^{\prime}$ (48)
$=\alpha_{k}(\tau_{k}-t)\exp{\int_{\tau_{k}-t}^{\tau_{k}}\gamma_{k}(s)}\mathrm{d}s.$ (49)